# FinCon: A Synthesized LLM Multi-Agent System with Conceptual Verbal Reinforcement for Enhanced Financial Decision Making

### v2

@@ 4.1.2 Risk-Control Component
4.1.3 FinCon Test Stage Workflow
@@
4.2 Modular Design of FinCon Agents
Here, we explain the modular design of FinCon agents. Inspired by the recent works of Park et al. [61] and Sumers et al. [71] on developing the cognitive structure of language agents for human-like behavior, agents in FinCon integrate four modules to support their necessary functionalities, along with a shared general configuration, as detailed in Figure 3:
General Configuration and Profiling Module. The general configuration defines the task types (e.g., stock trading, portfolio management) and specifies the trading targets, including background information like trading sectors and performance overviews. The profiling module outlines the roles and responsibilities of each agent in text. The textual content from these two parts is concatenated and used to query investment-related events stored in the agents’ memory databases.
Perception Module. This module defines how an agent interacts with the market, specifying what they perceive, receive, and send. These aspects of information vary among different functional agents, as illustrated in Figure 3.
@@ 4.2 Modular Design of FinCon Agents
Action Module Manager Agent:
1. Conduct: Trading actions.
@@ 1. Role assignment: You are an experienced trading manager in the investment firm …
1. Perceive: Investment insights from analyst agents & Risk signal and trajectory-level investment belief updates from the risk control component.
@@
2. Receive: Feedback from the manager agent.
@@ 1. Working:        - Observation    - Retrieval    - Distillation
Memory Module. The memory module contains three key components: working memory, procedural memory, and episodic memory. Similar to how humans process events in their working memory [72], FinCon agents utilize a working memory component to perform a series of operations, including observation, distillation, and refinement on available memory events, tailored to the agents’ specific roles. Procedural memory and episodic memory are essential for recording historical actions, outcomes, and reflections during sequential decision-making. Procedural memory is created after each decision step within an episode, storing data as memory events. For a trading inquiry, top events are retrieved from procedural memory, which are ranked based on recency, relevancy, and importance, which is the simplified version based on Yu et al. [23] and detailed in Appendix 7.6. Different functional analyst agents possesses distinct procedural memory decay rates, reflecting the timeliness of various financial data sources, which is crucial for aligning multi-type data impacting a specific time point and facilitating informed decision-making. The manager agent augments the procedural memory of analyst agents by providing feedback through access counter. Procedural memory is maintained by both analyst and manager agents, each keeping different records, as shown in Figure 3. The episodic memory, exclusive to the manager agent, retains actions, PnL series of the previous episodes and conceptualized updated investment beliefs from the risk control component.
5 Experiments
Reasearch Questions (RQs). Our experiment addresses three major research questions:
  • $\diamond$
    RQ1: Does FinCon demonstrate robustness across multiple financial decision-making tasks? We evaluate its performance on two challenging tasks: single-asset trading and portfolio management.
  • $\diamond$
    RQ2: Is the within-trajectory risk control mechanism in FinCon’s risk control component effective in maintaining superior decision-making performance?
  • $\diamond$
    RQ3: Is the over-trajectory risk control mechanism in FinCon’s risk control component effective for timely updating a manager agent’s beliefs and enhancing trading performance?
5.1 Experimental Setup
Multi-Modal Datasets. We collect a multi-modal dataset from real-world financial information to construct a representation of the market environment. It is composed of stock price information, daily financial news, company filing reports, and ECC audio recordings ranging from August 1, 2020, to August 14, 2023, as detailed in Appendix. 7.1. Each data source features distinct timeliness and is assigned to separate analyst agents to process. Typically, the time-sensitivity of annual filing reports (Form 10K’s) is more persistent, that of quarterly filing reports (Form 10Q’s) and ECC are secondary, and the daily financial news is the most instantaneous.
@@ 5.1 Experimental Setup
Evaluation Metrics. We evaluate the performance of FinCon and compare it with other state-of-the-art (SOTA) LLM-based and DRL-based agent systems using three standard financial performance metrics: Cumulative Return (CR%) [73], the Sharpe Ratio (SR) [74] and Max Drawdown (MDD%) [75]. Detailed definitions can be found in Appendix 7.3.
@@
Comparative Methods. For the single stock trading task, we assess FinCon’s performance against seven algorithmic agents and the Buy-and-Hold (B & H) strategy, a widely accepted baseline. These include three DRL-based agents from the FinRL framework [76]: A2C, PPO, and DQN, and four SOTA LLM-based agents: Generative Agent by Park et al. [11], FinGPT [77], FinMem [23], and FinAgent [24]. For the portfolio management task, FinCon competes against the classical Markowitz MV portfolio selection strategy [1], the RL-based FinRL-A2C agent [76] and the B & H strategy that is to hold an equal-weighted position in all trading assets (so-called equal-weighted ETF). Our focus on classical portfolio strategy, RL methods and B & H methods is due to the limited availability of mature LLM agents for portfolio management. The detailed experiment parameter configuration is detailed articulated in Appendix. 7.7.
@@
Implementation Details. In our experiments, all large language model (LLM)-based agent systems employ GPT-4-Turbo as their backbone algorithm, with the temperature parameter set at 0.3 to balance response consistency and reasoning creativity. FinCon is trained on financial data from January 3, 2022, to October 4, 2022, and is tested with data from October 5, 2022, to June 10, 2023. Given that deep reinforcement learning (DRL) algorithms require extensive training data for convergence, we extend the DRL agents’ training period to approximately five years, from January 1, 2018, to October 14, 2022, ensuring a fair comparison. The testing period remains consistent across all models. The performance metrics are reported for the test trajectory with the median CR and SR from five repeated epochs. (If the median CR and SR do not belong to the same epoch, the performance is based on the trajectory with the median CR.)
5.2 Main Results
In response to RQ1, we conduct analyses on two types of financial decision-making tasks: single-asset trading and portfolio management. The capability of FinCon to handle such sequentially complex decisions is thoroughly evaluated in the subsequent sections.
5.2.1 Single Asset Trading Task
In the first task, we evaluate the performance of FinCon against other leading algorithmic trading models by trading six different stocks. As shown in two tables above, FinCon significantly outperforms all the LLM-based and DRL-based approaches in terms of two primary metrics – cumulative returns and Sharpe Ratios. The FinCon’s Max Drawdown are also mostly among one of the lowest across all the trading assets, indicating a well-managed level of investment risk while still achieving the highest investment returns.
Overall, even with extended training periods, DRL-based models generally underperform. Specifically, the A2C algorithm significantly trails behind other algorithmic trading agents. Here, the training periods for Nio Inc (NIO) and Coinbase Global Inc (COIN) needs to be clarified. NIO completed its IPOs in September, 2018, so its training period is slightly shorter than other tickers. But the DRL algorithms for NIO still achieved convergence.The IPO for Coinbase Global, Inc (COIN) was completed in April 2021. With such limited trading data available, DRL algorithms struggle to converge, highlighting a significant limitation for DRL agents in automated trading of recent IPOs. Consequently, our analysis for COIN includes comparisons among FinCon, LLM-based trading agents, and the buy-and-hold strategy. In this analysis, FinCon demonstrates substantial performance advantages, achieving a cumulative return of over $57\%$ and a Sharpe ratio of $0.825$. We also note that LLM-based agents, capable of utilizing diverse data types and requiring minimal training, effectively address the challenges that DRL algorithms face.
Categories Models TSLA AMZN NIO
[…]
Market B&H 6.425 0.145 58.150 1.914 0.067 34.317 -77.210 -1.449 63.975
Our Model FinCon 82.871 1.972 29.727 24.964 0.906 25.889 17.461 0.335 40.647
LLM-based GA 16.535 0.391 54.131 -5.515 -0.195 37.213 -3.176 -1.574 3.155
FinGPT 1.549 0.044 42.400 -29.811 -1.805 -29.671 -4.959 -0.121 37.344
FinMem 34.624 1.552 15.674 -18.126 -0.776 36.825 -48.437 -1.180 64.144
FinAgent 11.960 0.271 55.734 -24.704 -1.496 33.151 0.933 0.051 19.181
DRL-based A2C -35.644 -0.805 61.502 -12.676 -0.447 37.179 -91.910 -1.728 68.911
PPO 1.409 0.032 49.740 3.863 0.137 28.085 -72.119 -1.352 62.093
DQN -1.296 -0.029 58.150 11.171 0.398 31.174 -35.419 -0.662 56.905
Categories Models APPL GOOG COIN
[…]
Market B&H 22.315 1.107 20.659 22.420 0.891 21.191 -21.756 -0.311 60.187
Our Model FinCon 27.352 1.597 15.266 25.077 1.0521 17.5299 57.045 0.825 42.679
LLM-based GA 5.694 0.372 14.161 -0.0151 -0.192 8.210 19.271 0.277 67.532
FinGPT 20.321 1.161 16.759 0.207 0.822 21.191 -99.553 -1.807 74.967
FinMem 12.396 0.994 11.268 0.311 0.018 21.503 0.811 0.017 50.390
FinAgent 20.757 1.041 19.896 -7.440 -1.024 10.360 -5.971 -0.106 56.882
DRL-based A2C 13.781 0.683 14.226 8.562 0.340 21.191 - - -
PPO 14.041 0.704 22.785 2.434 -0.097 25.202 - - -
DQN 21.125 1.048 16.131 20.690 0.822 21.191 - - -
Table 2: Comparison of key performance metrics during the testing period for the single-asset trading tasks involving six stocks, between FinCon and other algorithmic agents. Note that the highest and second highest CRs and SRs have been tested and found statistically significant using the Wilcoxon signed-rank test. The highest CRs and SRs are highlighted in red, while the second highest are marked in blue.
Moreover, in alignment with market trends, FinCon consistently exhibits superior sequential decision-making quality compared to other LLM-based financial agents, irrespective of whether the market conditions are bullish (e.g., GOOG), bearish (e.g., NIO), or mixed (e.g., TSLA). We attribute this performance to its high-quality distillation process through a synthesized multi-agent collaboration mechanism and its dual-functional risk control design, which collectively position FinCon as a leader in this space. Compared to FinCon, FinGPT primarily bases its trading decisions on the LLMs’ ability to interpret sentiment in financial information. This approach does not fully capitalize on the LLMs’ capacity to synthesize nuanced investment-related textual insights with numerical financial indicators. GA and FinMem, on the other hand, utilize single-agent frameworks that do not incorporate sophisticated information distillation processes or a diverse set of tools. This lack of complexity places considerable demand on the agent to process and reason through multi-sourced information, particularly when faced with large and varied data modalities. Additionally, they employ a static or minimal investment belief system. As a result, their ability to filter noise from mixed market data is relatively weak. As illustrated in Figure 7 (a) & (b), this limitation leads them to consistently maintain a lower position and exhibit hesitation regarding ‘buy’ or ‘sell’ decisions, ultimately resulting in suboptimal performance. FinCon overcomes these limitations through its innovative multi-agent synthesis, enabling it to deliver better outcomes. While FinAgent positions itself as a foundational framework capable of utilizing multi-modal data, its performance is robust with images and tabular data but does not exhibit a competitive edge when integrating audio data, such as ECC audio, which is crucial in real-world trading. Additionally, its decision-relevant memory retrieval process is based on similarity, leading the agent to make current decisions based on outdated memories, often resulting in errors. In contrast, the memory structure of FinCon is designed to account for the varying timeliness of multi-sourced financial data, ultimately delivering significantly superior decision quality.

### v3

@@ 3.1.2 Risk-Control Component
3.2 Modular Design of FinCon Agents
Here, we explain the modular design of FinCon agents. Inspired by the recent works of Park et al. [44] and Sumers et al. [45] on developing the cognitive structure of language agents for human-like behavior, agents in FinCon integrate four modules to support their necessary functionalities, along with a shared general configuration, as detailed in Appendix A.4:
General Configuration and Profiling Module. This module defines task types (e.g., stock trading, portfolio management) and specifies trading targets, including sector and performance details. The profiling module outlines each agent’s roles and responsibilities. The concatenated textual content from these parts is used to query investment-related events from the agents’ memory databases. Perception Module. This module defines how each agent interacts with the market, specifying the information they perceive, receive, and communicate, with interactions tailored to each agent’s role. In detail, it converts raw market data, feedback from other agents, and information retrieved from the memory module into formats compatible with large language models, enabling them to process these inputs effectively. Memory Module. The memory module comprises three key components: working memory, procedural memory, and episodic memory. Much like how humans process events in their working memory [46], FinCon agents leverage their working memory to perform a range of tasks, including observation, distillation, and refinement of available memory events, all tailored to the specific roles of the agents. Procedural memory and episodic memory are critical for recording historical actions, outcomes, and reflections during sequential decision-making. Procedural memory is generated after each decision step within an episode, storing data as memory events. For trading inquiries, top events are retrieved from procedural memory and ranked based on recency, relevance, and importance, following a simplified version of the method proposed by Yu et al. [32], with further details provided in Appendix A.13. Each functional analyst agent has distinct procedural memory decay rates, reflecting the timeliness of various financial data sources, which is crucial for aligning multi-type data influencing specific time points and supporting informed decision-making. The manager agent enhances the procedural memory of analyst agents by providing feedback through an access counter. Both analyst and manager agents maintain procedural memory, but they keep different records, as illustrated in Appendix A.4. Episodic memory, exclusive to the manager agent, stores actions, PnL series from previous episodes, and updated conceptual investment beliefs from the risk control component.
4 Experiments
Our experiment answers the key research questions (RQs): RQ1: Does FinCon demonstrate robustness across multiple financial decision-making tasks, especially single-asset trading and portfolio management? RQ2: Is the within-episode risk control mechanism in FinCon effective in maintaining superior decision-making performance? RQ3: Is the over-episode risk control mechanism in FinCon effective in timely updating the manager agent’s beliefs to enhance trading performance?
4.1 Experimental Setup
(i) Multi-Modal Datasets. We construct a market environment representation using real-world financial data, including stock prices, daily news, company filings (Form 10-Q, Form 10-K, etc.), and ECC audio from January 3, 2022, to June 10, 2023, as detailed in Appendix. A.8. Each data source is assigned to specific analyst agents based on its timeliness. (ii) Evaluation Metrics. We evaluate FinCon and other state-of-the-art (SOTA) agents using metrics such as Cumulative Return (CR%), Sharpe Ratio (SR), and Max Drawdown (MDD%). CR and SR are prioritized because they provide comprehensive insights into overall performance and risk-adjusted returns, essential for informed investment decisions. In contrast, MDD focuses on evaluating the potential for significant losses, making it a secondary consideration in this context. Details are provided in Appendix A.10. (iii) Comparative Methods. For single-stock trading, we compare FinCon with DRL agents (A2C, PPO, DQN) and LLM-based agents (Generative Agent (GA), FinGPT, FinMem, FinAgent) as well as the Buy-and-Hold (B & H) strategy. For portfolio management, we compare FinCon with Markowitz MV, FinRL-A2C, and Equal-Weighted ETF strategy, with further details provided in Appendix A.12. The detailed experiment parameter configurations of the above agent systems are articulated in Appendix. A.14. (iv) Implementation Details. All LLM-based agents use GPT-4-Turbo, with temperature set at 0.3. FinCon is trained from January 3, 2022, to October 4, 2022, and tested from October 5, 2022, to June 10, 2023. DRL agents are trained over the period from January 1, 2018, to October 4, 2022, to ensure that there is sufficient data available for model convergence. Performance is based on the median CR and SR from five repeated epochs. For a more detailed explanation of the experimental setup, please refer to the Appendix A.5.
4.2 Main Results
In response to RQ1, we analyze FinCon’s performance on two types of financial decision-making tasks: single-asset trading and portfolio management. The system’s ability to manage these sequentially complex decisions is thoroughly evaluated in the following sections.
4.2.1 Single Asset Trading Task
In this task, we evaluate FinCon’s performance against other leading algorithmic trading models by trading eight different stocks. As presented in the tables above, FinCon significantly outperforms both LLM-based and DRL-based approaches in terms of CRs and SRs. Additionally, FinCon achieves one of the lowest MDD values across most trading assets, demonstrating effective risk management while still delivering the highest investment returns. For detailed performance comparisons across all models and metrics, refer to Table 1.
Overall, even with extended training periods, DRL-based models tend to underperform, with the A2C algorithm lagging significantly behind other agents in general. Notably, the training periods for Nio Inc. (NIO) and Coinbase Global Inc. (COIN) require clarification. NIO, which completed its IPO in September 2018, has a slightly shorter training period than other tickers, yet the DRL algorithms for NIO still achieved convergence. In contrast, Coinbase Global Inc. (COIN), which completed its IPO in April 2021, presented a more significant challenge due to the limited available trading data, causing DRL algorithms to struggle with convergence. This limitation underscores a major drawback for DRL agents when trading recently listed IPOs. Consequently, our analysis of COIN focuses on comparisons between FinCon, LLM-based agents, and the buy-and-hold (B & H) strategy. In this context, FinCon demonstrates a clear advantage, achieving a cumulative return of over 57% and a Sharpe ratio of 0.825. Furthermore, LLM-based agents, which can leverage diverse data types and require minimal training, effectively mitigate the challenges faced by DRL algorithms.
Categories Models TSLA AMZN NIO MSFT
[…]
Market B&H 6.425 0.145 58.150 2.030 0.072 34.241 -77.210 -1.449 63.975 27.856 1.230 15.010
Our Model FinCon 82.871 1.972 29.727 24.848 0.904 25.889 17.461 0.335 40.647 31.625 1.538 15.010
LLM-based GA 16.535 0.391 54.131 -5.631 -0.199 37.213 -3.176 -1.574 3.155 -31.821 -1.414 39.808
FinGPT 1.549 0.044 42.400 -29.811 -1.810 29.671 -4.959 -0.121 37.344 21.535 1.315 16.503
FinMem 34.624 1.552 15.674 -18.011 -0.773 36.825 -48.437 -1.180 64.144 -22.036 -1.247 29.435
FinAgent 11.960 0.271 55.734 -24.588 -1.493 33.074 0.933 0.051 19.181 -27.534 -1.247 39.544
DRL-based A2C -35.644 -0.805 61.502 -12.560 -0.444 37.106 -91.910 -1.728 68.911 21.397 0.962 21.458
PPO 1.409 0.032 49.740 3.863 0.138 28.085 -72.119 -1.352 62.093 -4.761 -0.214 30.950
DQN -1.296 -0.029 58.150 11.171 0.398 31.174 -35.419 -0.662 56.905 27.021 1.216 21.458
Categories Models AAPL GOOG NFLX COIN
[…]
Market B&H 22.315 1.107 20.659 22.420 0.891 21.191 57.338 1.794 20.926 -21.756 -0.311 60.187
Our Model FinCon 27.352 1.597 15.266 25.077 1.052 17.530 69.239 2.370 20.792 57.045 0.825 42.679
LLM-based GA 5.694 0.372 14.161 -1.515 -0.192 8.210 41.770 1.485 20.926 19.271 0.277 67.532
FinGPT 20.321 1.161 16.759 0.242 0.011 26.984 11.925 0.472 20.201 -99.553 -1.807 74.967
FinMem 12.397 0.994 11.268 0.311 0.018 21.503 -10.306 -0.478 27.692 0.811 0.017 50.390
FinAgent 20.757 1.041 19.896 -7.440 -1.024 10.360 61.303 1.960 20.926 -5.971 -0.106 56.882
DRL-based A2C 13.781 0.683 14.226 8.562 0.340 21.191 -8.176 -0.258 49.579 - - -
PPO 14.041 0.704 22.785 2.434 0.097 25.202 -33.144 -1.049 33.377 - - -
DQN 21.125 1.048 16.131 20.690 0.822 21.191 21.753 0.687 39.733 - - -
Table 2: Comparison of key performance metrics during the testing period for the single-asset trading tasks involving eight stocks, between FinCon and other algorithmic agents. Note that the highest and second highest CRs and SRs have been tested and found statistically significant using the Wilcoxon signed-rank test. The highest CRs and SRs are highlighted in red, while the second highest are marked in blue.
In alignment with market trends, FinCon consistently exhibits superior decision-making quality compared to other LLM-based agents, regardless of market conditions—whether bullish (e.g., GOOG, MSFT), bearish (e.g., NIO), or mixed (e.g., TSLA). We attribute this performance to its high-quality distillation of information through a synthesized multi-agent collaboration mechanism, combined with its dual-level risk control design, positioning FinCon as a leader in the space. By contrast, FinGPT primarily relies on sentiment analysis of financial information, failing to fully exploit the potential of LLMs to integrate nuanced textual insights with numerical financial indicators. Similarly, GA and FinMem use single-agent frameworks without sophisticated information distillation processes or a diverse toolset, placing heavy cognitive demand on the agent to process multi-source information, especially when dealing with large and varied data modalities. Moreover, their static or minimal investment belief systems result in weak filtering of market noise. As illustrated in Figure 7 (a) & (b) of Appendix A.7.2, this limitation leads these models to consistently hold lower positions and hesitate between ‘buy’ or ‘sell’ decisions, ultimately resulting in suboptimal performance.
FinCon overcomes these challenges through its innovative multi-agent synthesis, enabling it to deliver superior outcomes. Although FinAgent performs well when integrating images and tabular data, it struggles to remain competitive when incorporating audio data, such as ECC recordings, which are critical in real-world trading. Additionally, FinAgent relies on similarity-based memory retrieval, which can lead to decisions based on outdated information, often resulting in errors. In contrast, FinCon’s memory structure accounts for the varying timeliness of multi-source financial data, significantly enhancing decision quality and overall performance.
4.2.2 Portfolio Management Task
In this task, we compare FinCon’s performance with the Markowitz Mean-Variance (MV) portfolio [47] and FinRL [48] in managing two small portfolios: Portfolio 1 (TSLA, MSFT, and PFE) and Portfolio 2 (AMZN, GM, and LLY). These assets were selected by the stock selection agent from a pool of 42 stocks, each with sufficient news data (over 800 news articles during the combined training and testing periods), as illustrated in Figure 9 in Appendix A.9. The training and testing periods, the backbone model and the parameter settings are consistent with those used in the single-asset trading task. For the Markowitz MV portfolio, we estimate the covariance matrix and expected returns using the same training data. In the case of FinRL, we use five years of training data prior to the test period. As detailed in Table 3 and Figure 3, our results show that FinCon outperforms both the Markowitz MV portfolio and FinRL as well as the market baseline – Equal-Weighted ETF, achieving significantly higher CRs and SRs, as well as MDDs.
However, managing multi-asset portfolios introduces more complexity, leading to a higher likelihood of hallucination compared to single-asset trading. This is due to the increased input length and complexity involved in multi-asset decision-making. While FinCon mitigates this issue by distributing tasks across specialized agents that focus on critical investment insights, it occasionally generates incorrect information, such as non-existent indices of memory events. Handling multi-asset decision-making requires sophisticated logic and substantial market information, which poses a significant challenge for LLMs when processing extended contexts. This complexity has left portfolio management relatively unexplored in previous language agent studies. Nonetheless, FinCon demonstrates considerable potential by constructing agent systems that can tackle complex financial tasks through effective resource optimization, even when managing relatively compact portfolios.
[…]
Table 3: Key performance metrics comparison among all portfolio management strategies of Portfolio 1 & 2. FinCon leads all performance metrics.
Figure 3: Portfolio values of Portfolio 1 & 2 changes over time for all the strategies. The computation of portfolio value refers to Equation 7 in Appendix A.10.
4.3 Ablation Studies
In response to RQ2 and RQ3, we conduct a comprehensive evaluation of our unique risk control component through two ablation studies. Both studies maintain consistency with the training and testing periods used in the main experiments. The first study examines the effectiveness of the within-episode risk control mechanism, which leverages Conditional Value at Risk (CVaR) to manage risk in real-time, as detailed in Table 4. Comparisons on primary metrics illustrate that the success of utilizing CVaR for within-episode risk control is evident in both bullish and bearish market environments in the single asset trading case. Moreover, in portfolio trading with mixed price trends, our within-episode risk control mechanism performs robustly by monitoring the entire portfolio’s value fluctuations. The second study focuses on the over-episode risk control mechanism, demonstrating its critical role in updating the trading manager agent’s beliefs to provide a more comprehensive understanding of current trading conditions, as articulated in Table 5. The markedly improved CRs and SRs in both decision-making scenarios underscore the effectiveness of using CVRF to update investment beliefs episodically, guiding the agent towards more profitable investment strategies. Additionally, FinCon demonstrates significant learning gains, achieving these results after only four training episodes—substantially fewer than what is typically required by traditional RL algorithmic trading agents. More visualizations and analysis are provided in the Appendix A.7.
[…]
Table 5: Key metrics FinCon with vs. without implementing belief updates for over-episode risk control. The performance of FinCon with the implementation of CVRF won a leading performance in both single-asset trading and portfolio management tasks.
5 Conclusion
In this paper, we present FinCon, a novel LLM-based multi-agent framework for financial decision-making tasks, including single stock trading and portfolio management. Central to FinCon is the Synthesized Manager-Analyst hierarchical communication structure and a dual-level risk control component. This communication method channels financial data from multiple sources to specialized analyst agents, who distill it into key investment insights. The manager agent then synthesizes these insights for decision-making. Our experimental evaluations demonstrate the efficacy of our risk control mechanism in mitigating investment risks and enhancing trading performance. Additionally, the streamlined communication structure reduces overhead. The dual-level risk control component introduces a novel approach to defining agent personas, enabling dynamic updates of risk and market beliefs within agent communication. A valuable future research direction would be to scale FinCon’s framework to manage large-sized portfolios comprising tens of assets, while maintaining the impressive decision-making quality demonstrated with smaller portfolios. Given the LLM’s input length constraint, a critical challenge lies in striking an optimal balance between information conciseness through agent distillation and potential performance deterioration when extending the current context window. Addressing this will be essential for ensuring quality-assured outcomes.
Appendix A Appendix
A.1 Related Work
LLM Agents for Financial Decision Making. There are considerable efforts towards developing general-purpose LLM agent for sequential decision-making [49, 50], and such type of tasks often involve episodic interactions with environment and verbal reflections for action refinement, such as coding competition [51, 52], software development [53, 23], game-playing [54, 55]. Furthermore, researchers have started to exploit how LLM agents can perform better in harder decision-making tasks from finance [56, 57, 58, 59, 60], in which there are more volatile environments, leading to that the numerous unpredictable elements can obscure an agent’s ability to reflect accurately on the reasons for poor decision outcomes. FinMem [32] enhances single stock trading performance by embedding memory modules with LLM agent for reflection-refinement, and FinAgent [33] improved trading profits via using external quantitative tool to fight against volatile environment.
Multi-Agent System and Communication Structures. In traditional multi-agent systems [61, 62], the way for agents’ communication is pre-determined, like sharing data or state observations [63, 64, 65, 66, 67, 68]. The emergence of large language model brings flexibility for human-understandable communications [69, 20, 23, 70], so some work tries to elevate decision-making ability of LLM-based multi-agent system by letting agents engage in discussions [71, 21] or debates [72, 73]. The similar peer-communication strategy was as well utilized by the multi-agent system for financial tasks [74, 75, 76]. However, such approach are not optimal for unified-goal financial tasks that prioritize profits [77], because they suffer from potentially ambiguous optimization objectives and are unable to control the unnecessary communication costs [78].
Prompt Optimization and Verbal Reinforcement. To enhance the reasoning or decision-making of LLM agents, many prompt optimization techniques have been proposed, like ReAct [79], Chain of Thought (CoT) [80], Tree of Thoughts (ToT) [81], ART [14], intended for that LLM agents can automatically generate intermediate reasoning steps as an iterative program. In addition, to make LLM agents make decisions like humans and generate more understandable reasoning texts, some researchers recommend incorporating cognitive structures [82, 83, 44, 84]. Inspired by these previous work and DRL algorithms [85, 86, 87, 67, 88], verbal reinforcement [29, 30, 89, 24] was developed for LLM agents such that they can update actions based on iterative self-reflection while integrating additional LLM as a prompt optimizer [27, 28].
A.2 Textual Gradient-Descent
A.3 FinCon Testing Stage Workflow
@@
A.4 Figure of Modular Design of Agents in FinCon
@@ 1. Role assignment: You are an experienced trading manager in the investment firm …
1. Perceive: Investment insights from analyst agents; daily risk alert and episode-level investment belief updates from the risk control component.
@@
2. Send: Relevant market insights to manager agent. 3. Receive: Feedback from the manager agent.
@@ 1. Working:        - Observation    - Retrieval    - Distillation
Action Module Manager Agent:
1. Conduct: Trading actions.
@@
A.5 Experimental Setup
Multi-Modal Datasets. We collect a comprehensive multi-modal dataset to simulate a realistic market environment. This dataset includes stock price data, daily financial news, company filing reports (10K and 10Q), and ECC (Earnings Call Conference) audio recordings spanning from January 3, 2022, to June 10, 2023. Each data source is assigned to specific analyst agents based on the timeliness of the information. For example, annual filings (10K) exhibit longer-term persistence, quarterly filings (10Q) and ECC data have medium-term relevance, and daily financial news provides the most immediate information.
@@ A.5 Experimental Setup
Evaluation Metrics. We evaluate FinCon and benchmark it against other state-of-the-art LLM-based and DRL-based agent systems using three key financial performance metrics: Cumulative Return (CR%) [90], Sharpe Ratio (SR) [91], and Max Drawdown (MDD%) [92]. These metrics help quantify each model’s profitability, risk-adjusted returns, and risk management performance, respectively.
@@
Comparative Methods. In the single-stock trading task, we compare FinCon against seven algorithmic agents and the widely accepted Buy-and-Hold (B & H) baseline. The three DRL-based agents—A2C, PPO, and DQN—are from the FinRL framework [48], while the four state-of-the-art LLM-based agents include Generative Agent [20], FinGPT [93], FinMem [32], and FinAgent [33]. For portfolio management, we benchmark FinCon against the classical Markowitz MV portfolio selection strategy [1], the RL-based FinRL-A2C agent [48], and the B & H strategy, which holds an equal-weighted position across all assets (equal-weighted ETF). Our focus on classical, RL-based, and B & H methods is due to the current lack of mature LLM-based agents for portfolio management tasks.
@@
Implementation Details. In our experiments, all LLM-based agent systems, including FinCon, use GPT-4-Turbo as the backbone model, with the temperature parameter set at 0.3 to balance response consistency with creative reasoning. FinCon is trained on financial data from January 3, 2022, to October 4, 2022, and tested on data from October 5, 2022, to June 10, 2023. Since deep reinforcement learning (DRL) agents require extensive data for convergence, their training period is extended to nearly five years (January 1, 2018, to October 4, 2022) to ensure fair comparison. The testing period remains the same across all models. The final performance metrics are based on the test trajectory with the median CR and SR values from five repeated epochs. If the median CR and SR occur in different epochs, performance is assessed based on the trajectory with the median CR value.
A.6 Single Stock Trading Result Graphs